OpenStudy (anonymous):

Find the perimeter of a rectangle? Given: Diagonal=7is square route of 5 area= to 98

OpenStudy (loser66):

Do you know the formula of area of a rectangle?

OpenStudy (anonymous):

Area is = LXW

OpenStudy (anonymous):

and the perimeter is 2b+2h

OpenStudy (anonymous):

If Im not mistaken I should find first the Length and the Width then substitute the values then solve for the perimeter now

OpenStudy (loser66):



OpenStudy (loser66):

now, we need W and L right?

OpenStudy (anonymous):

thats it !

OpenStudy (anonymous):

Should i use the Pythagorean theorem

OpenStudy (loser66):

do you know pythagorean theorem?

OpenStudy (anonymous):

c^2=a^2+b^2

OpenStudy (loser66):

hey!! should use the same notation, in this case it is??

OpenStudy (anonymous):

I can get the diagonal \[7\sqrt{5}\] =a^2 + b^2

OpenStudy (loser66):

I realize that you know everything. However, you will mess it up by using different notation on 1 problem

OpenStudy (anonymous):

but im a little bit confuse. i tried this formula look.

OpenStudy (loser66):

W*L= 98 \((7\sqrt5)^2= W^2+L^2\)

OpenStudy (anonymous):